OpenStudy (anonymous):

Find the difference. (-6x + 2y -3) - (2x - 4y - 7)

OpenStudy (anonymous):

-4x - 2y - 4 -8x - 2y - 10 -8x + 6y + 4

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

No ! OK ! Again !:) -6x-2x=-8x 2y+4y=10y 7-3=4

OpenStudy (anonymous):

Noonoo Its 6y not 10y. ! it s C .!

OpenStudy (anonymous):

@charlieberzak :OK ???

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

OK :))